# A Statistical Test for Comparing the Linkage and Admixture Model Based on Central Limit Theorems

### v1

@@ 5 Proofs of the Main Results
We first prove that the MLE is asymptotically unique. Based on this, we prove consistency and central limit results for the MLE in the Linkage Model. For all proofs, we assume $C=1$ to simplify the notation.
@@ 5.1 Unique MLE
In this section, we use that the limit
$\ell((X_{1},...,X_{M}),(q,r))\xrightarrow[]{M\to\infty}\ell(q,r)$
exists almost surely. We will prove this in section 5.2.
Proving the uniqueness of a MLE for HMM has been already considered extensively for a finite state space, see e.g. (Finesso, 1990; Gilbert, 1959; Blackwell and Koopmans, 1957; Petrie, 1969). In this section, we aim to prove the asymptotic uniqueness of the MLE in the Linkage Model, i.e. we aim to prove Theorem 4.
Theorem 4.
@@
Therefore, we first prove that the parameters are identifiable based on a finite number of observations. Based on this, we prove that if $\ell(q^{0},r^{0})\neq\ell(q,r)$, then the measures $\mathbb{P}_{q,r}|_{(X_{m})_{m=1,2,...}}$ and $\mathbb{P}_{q^{0},r^{0}}|_{(X_{m})_{m=1,2,...}}$ are mutually different. Based on this, we can then conclude the claim.
We write $(q,r)\sim(q^{\prime},r^{\prime})$, if there exists an invariant distribution $\pi,\pi^{\prime}$ such that $\mathbb{P}_{(q,r)}^{X}=\mathbb{P}_{(q^{\prime},r^{\prime})}^{X}$ holds. Therefore, we only consider the case $K=2$ as the proof involves naming conditions that yield to the full rank of the emission matrix.
@@
Let Assumption 3.1 hold and let $K=2.$ If either the allele frequencies $p_{\cdot,i},...,p_{\cdot,i+s}$ are linearly independent and $s\geq 4$ or $s=3$ and $q\neq p_{1,2}/(p_{1,2}+p_{2,2})$ holds and in addition $p_{1,m}\neq p_{2,m}\forall m\in\{i,...,i+s\}$, the parameters $(q,r)$ are identifiable based on the random variables $X_{i},...,X_{i+s}.$
@@
and $M_{1},M_{2},M_{3}]$ is a three-dimensional array for which the element $(x_{1},x_{2},x_{3})$ equals $M_{1},M_{2},M_{3}]_{x_{1},x_{2},x_{3}}.$ It holds
$\mathbb{P}(X_{1},X_{2},X_{3})$ $=\sum_{z_{1},z_{2},z_{3}}\pi_{z_{1}}A_{1}(z_{1},z_{2})A_{2}(z_{2},z_{3})B_{2}(X_{2},z_{2})B_{3}(X_{3},z_{3})B_{1}(X_{1},z_{1})$
$=[\underbrace{\pi B_{1}A_{1}}_{=:M_{1}},B_{2},\underbrace{A_{2}B_{3}}_{=:M_{2}}].$
@@
    prove that the entries of $M_{1},M_{2},B_{2}$ are identifiable from the marginal density of $X_{1},X_{2},X_{3}$. Therefore, we use theory by Allman et al. (2009); Kruskal (1977).
@@
Hence, $B_{2}$ has the full rang if and only if $q_{1}\neq\frac{p_{1,2}}{p_{1,2}+p_{2,2}}.$ This is the reason, why we need the condition about the linear independence of the allele frequencies (since we can apply Kruskal’s theorem to two different time points within the interval). We have
@@
$a$ $=(1-q_{1}p_{1,1})\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right)+\left(1-(1-q_{1})p_{2,1}\right)(1-e^{-rd_{1}})(1-q_{1})$
@@
$b$ $=q_{1}p_{1,1}\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right)+p_{2,1}(1-q_{1})^{2}(1-e^{-rd_{1}}),$
$c$ $=(1-q_{1}p_{1,1})(1-e^{-rd_{1}})(1-q_{1})+\left(1-(1-q_{1})p_{2,1}\right)\left((1-e^{-rd_{1}})(1-q_{1})+e^{-rd_{1}}\right),$
$d$ $=q_{1}p_{1,1}(1-e^{-rd_{1}})(1-q_{1})+p_{2,1}(1-q_{1})\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right).$
@@
$\det(M_{1})=e^{-rd_{1}}(1-e^{-rd_{1}})q_{1}(1-q_{1})(p_{1,1}+p_{2,1}-2p_{1,1}p_{2,1})=0$
$\Leftrightarrow(r=\infty)\lor(r=0)\lor(q_{1}\in\{0,1\})\lor(p_{1,1}=p_{2,1}=0)\lor(p_{2,1}=p_{1,1}=1).$
Here, we excluded every case except the case $r=\infty.$ This case leads to the Admixture Model which has already been discussed by Heinzel (2025). We additionally calculate
$\mathbb{P}(X_{k-2}=x\mid Z_{k-1}=z)$ $=\sum_{z^{\prime}}\mathbb{P}(X_{k-2}=x,Z_{k-2}=z^{\prime}\mid Z_{k-1}=z)$
$\propto\sum_{z^{\prime}}\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=z^{\prime})\cdot\mathbb{P}(Z_{k-2}=z^{\prime}\mid Z_{k-1}=z)$
$\propto\sum_{z^{\prime}}\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=z^{\prime})\cdot\mathbb{P}(Z_{k-1}=z^{\prime}\mid Z_{k-2}=z)$
@@
$\propto(q_{1}p_{1,3})^{x}(1-q_{1}p_{1,3})^{1-x}\left(1_{z=0}e^{-rd_{2}}+q_{z}(1-e^{-rd})\right)$
$\hskip 28.45274pt+(q_{2}p_{2,3})^{x}(1-q_{2}p_{2,3})^{1-x}\left(1_{z=1}e^{-rd_{2}}+q_{z}(1-e^{-rd})\right).$
@@
$a_{2}$ $=(1-q_{1}p_{1,3})\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right)+(1-(1-q_{1})p_{2,3})q_{1}(1-e^{-rd_{2}})$
$b_{2}$ $=(q_{1}p_{1,3})\left(e^{-rd_{2}}+q_{1}(1-e^{-rd_{2}})\right)+((1-q_{1})p_{2,3})q_{1}(1-e^{-rd_{2}}),$
$c_{2}$ $=(1-q_{1}p_{1,3})(1-q_{1})(1-e^{-rd_{2}})+(1-(1-q_{1})p_{2,3})\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right),$
$d_{2}$ $=(q_{1}p_{1,3})(1-q_{1})(1-e^{-rd_{2}})+((1-q_{1})p_{2,3})\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right).$
Hence, it holds $rk(M_{2})=1\Leftrightarrow p_{2,3}=p_{1,3},q_{1}=0.5.$ Now, we come to step (ii), i.e. we have to ensure that the parameters are uniquely determined based on the entries. For the emission probabilities, this is a direct consequence of the assumption $p_{k,m}\neq p_{\ell,m}.$ Then, since $d_{m}>0,$ it is also immediate to infer the value $r$ from the other parameters.
Remark 5.2 (Identifiability in the Admixture Model for $K=2$).
Heinzel (2025) did not consider the identifiability for finitely many markers in the Admixture Model. For $K=2$, it holds
$\mathbb{P}(X_{m}=x)$ $=\langle q,p_{\cdot,m}\rangle^{x}(1-\langle q,p_{\cdot,m}\rangle)^{1-x}$
$=(q_{1}p_{1,m}+(1-q_{1})p_{2,m})^{x}(1-q_{1}p_{1,m}-(1-q_{1})p_{2,m})^{1-x}$
$=\begin{cases}p_{2,m}+q_{1}(p_{1,m}-p_{2,m})=:c_{m},x=1\\ 1-p_{2,m}+q_{1}(-p_{1,m}+p_{2,m}),x=0\end{cases}.$
This leads to
$q_{1}=\frac{c_{m}-p_{2,m}}{p_{1,m}-p_{2,m}},$
if $p_{1,m}=p_{2,m}.$ Hence, $q_{1}$ is uniquely determined if and only if $p_{1,m}=p_{2,m}.$
Remark 5.3.
@@
Now, we aim to infer (3) from the identifiability of the parameters. Therefore, we first prove a law of large numbers. Then, based on this, we conclude (3). This is done by using the following particularity of the hidden chain (Lemma 5.4) and that $(Z_{m})_{m=1,2,...}$ is uniformly ergodic (Lemma 5.6).
@@
Lemma 5.6 (Uniformly Ergodicity for the Linkage Model).
@@
This claim is derived from Theorem 3.3 in Saloff-Coste and Zúniga (2007). ∎
Now, we are ready to name an LLN for the Linkage Model.
@@
Let Assumption 3.1 hold and let $f_{i}$ be uniformly bounded by $f_{\infty}$. Then, it holds
$\mathbb{P}^{\nu}_{\theta}\left(\frac{1}{n}\left|\sum_{i=1,...,n}f_{i}(X_{i:i+s})-\pi(f_{i})\right|\geq\epsilon\right)\leq C\exp\left(-\frac{n}{8f_{\infty}}\right)$
for a constant $C>0$ and for any probability measure $\nu.$
Proof.
Without loss of generality, we assume $\mathbb{E}_{\pi}(f_{i}(X_{i:i+s}))=0$. From Lemma 5.6, we know that the Markov chain $(Z_{m})_{m=1,2,...}$ is uniformly ergodic. The proof is similar to Theorem 14 in Douc (2005): We first write
$\xi_{i,j}:=\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|Z_{0:i+j},X_{0:i+j})+\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|Z_{0:i+j-1},X_{0:i+j-1}).$
Then, since
$\sum_{j=0}^{s}\sum_{i=1}^{n}\xi_{i,j}$ $=\sum_{i=1}^{n}\sum_{j=0}^{s}\xi_{i,j}=\sum_{i=1}^{n}\xi_{i,s}-\xi_{i,0},$
it holds
$\sum_{i=1,...,n}f_{i}(X_{i:i+s})=\sum_{j=0}^{s}\sum_{i=1}^{n}\xi_{i,j}+\sum_{i=1}^{n}\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|Z_{0:i-1},X_{0:i-1}).$
Now, we can apply Azuma-Hoeffding inequality, which leads to
$\mathbb{P}^{\nu}\left(\bigg{|}\sum_{i=1}^{n}\xi_{i,j}\bigg{|}\geq t\right)\leq 2\exp\left(-\frac{t^{2}}{8nf_{\infty}}\right).$
Now, we note that $\sum_{i=1}^{n}\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|X_{0:i-1},Z_{0:i-1})=F(Z_{i-1})$ for a certain function $F.$ We now prove that
$\mathbb{P}\left(\bigg{|}\sum_{i=1}^{n}\mathbb{E}_{\nu}(f_{i}(X_{i:i+s})|X_{0:i-1},Z_{0:i-1})\bigg{|}\geq t\right)2\leq\exp\left(-\frac{t^{2}}{8nf_{\infty}}\right).$
This is Theorem 17 in Douc et al. (2011).
To prove this claim, we apply Etemadi’s inequality and Bernstein’s inequality to the $\xi_{k}$ as defined in Lemma 18 in Douc et al. (2011). Neither of them requires that the $\xi_{k}$ are identically distributed. Additionally, they used their Proposition 19 to prove the claim. Note, however, that in our case we can highly simplify the proof as we have a discrete Markov chain with bounded transition probabilities. Especially, we do not even need their $\check{X}_{n}$ as in our case, their $d_{n}$ equals 1 for all $n\in\mathbb{N}.$
Lemma 5.8.
Assumption 3.1 implies that there exists a uniformly bounded function $h_{i}$ and an integer $s\in\mathbb{N}$ such that
$(i)\quad\frac{1}{M}\sum_{i=1}^{M}h_{i}(X_{i},...,X_{i+s})$ $\xrightarrow[]{M\to\infty}1\quad\mathbb{P}_{q^{0},r^{0}}-a.s.,$
$(ii)\quad\frac{1}{M}\sum_{i=1}^{M}h_{i}(X_{i},...,X_{i+s})$ $\xrightarrow[]{M\to\infty}0\quad\mathbb{P}_{q,r}-a.s.$
@@
The proof is based on Lemma 6 in Douc et al. (2011). However, in our case $X$ is not a stationary sequence. We find an integer $s$ and bounded functions $h_{i},i=1,2,$..., such that $\mathbb{E}_{\pi}(h_{i}(X_{i},...,X_{i+s}))=1,\mathbb{E}_{\pi^{0}}(h_{i}(X_{i},...,X_{i+s}))=0$ according to the theorem from Hahn-Banach for $\pi\neq\pi^{0}$. This is a consequence of Lemma 5.1. More precisely, we choose
$h_{i}(X_{i},...,X_{i+s})=1-\frac{f_{i}(X_{i},...,X_{i+s})-1}{\mathbb{E}_{\pi^{0}}(f_{i}(X_{i},...,X_{i+s}))-1}$
with
$f_{i}(X_{i},...,X_{i+s}):=\frac{\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s})}{\mathbb{P}_{\pi}(X_{i},...,X_{i+s})}$
for $\pi\neq\pi^{0}.$ The case $\mathbb{E}_{\pi^{0}}(f_{i})=1$ cannot occur, which follows from Cauchy-Schwarz. However, the densities are uniformly bounded according to Assumption 3.1, i.e. we can also choose $h_{i}$ uniformly bounded.
Hence, we use Dolgopyat and Sarig (2023), Theorem 3.12, to prove (i).
For (ii), We define
$A_{M}:=\left\{X_{0:M}:\frac{1}{M}\sum_{i=1}^{M}h_{i}(X_{i},...,X_{i+s})\geq\frac{1}{2}\right\}.$
It holds
$\mathbb{P}_{q,r}^{\nu}(X_{0:M}\in A_{M})$
$\quad=\sum_{k=1}^{K}\mathbb{P}_{q,r}^{\nu}(X_{0:M}\in A_{M}|Z_{0}=k)\mathbb{P}_{q,r}^{\nu}(Z_{0}=k)$
$\quad\leq\max_{k=1,...,K}\sup_{\textup{supp }\mu}\mathbb{P}^{\mu}_{q,r}\left(\frac{1}{M}\sum_{i=1}^{M}h_{i}(X_{i},...,X_{i+s})\geq\frac{1}{2}\right)$
$\quad\leq c_{1}\exp(-M/c_{2})$
for some constant $c_{1},c_{2}>0.$ Here, we use Lemma 5.7 for the last inequality. ∎
Finally, we use Lemma 5.8 to prove Theorem 4.
@@
Since the number of observable states is finite, we immediately obtain this.
Now, we prove that $\mathbb{P}_{q^{0},r^{0}}\neq\mathbb{P}_{q,r}$ implies that $\mathbb{P}_{q^{0},r^{0}},\mathbb{P}_{q,r}$ are mutually singular. This is Lemma 5.8. Hence, we can then infer the claim analogously to van Handel (2008) in the proof of Theorem 7.13. ∎

### v2

@@ 5 Proofs of the Main Results
We first prove that the MLE $\left(\hat{Q}^{M},\hat{R}^{M}\right):=\left(\hat{Q}^{1,M},\hat{R}^{1,M}\right)$ is asymptotically unique. Based on this, we prove consistency and central limit results for the MLE in the Linkage Model. For all proofs, we assume $C=1$ to simplify the notation.
@@ 5.1 Unique MLE
Proving the uniqueness of a MLE for HMM has been already considered extensively for a finite state space, see e.g. (Finesso, 1990; Gilbert, 1959; Blackwell and Koopmans, 1957; Petrie, 1969). In this section, we aim to prove the asymptotic uniqueness of the MLE in the Linkage Model, i.e. we aim to prove Theorem 4. Therefore, we use that the limit
$\ell((X_{1},...,X_{M}),(q,r))\xrightarrow[]{M\to\infty}\ell(q,r)$ (3)
exists $\mathbb{P}$-almost surely. We will prove this in section 5.2.
Theorem 4 (Unique Maximum Likelihood Estimator).
@@
To prove Theorem 4, we first prove that the parameters are identifiable based on a finite number of observations (Lemma 5.1). Based on this, we prove that if $\ell(q^{0},r^{0})\neq\ell(q,r)$, then the measures $\mathbb{P}_{q,r}|_{(X_{m})_{m=1,2,...}}$ and $\mathbb{P}_{q^{0},r^{0}}|_{(X_{m})_{m=1,2,...}}$ are mutually different. Then, we can conclude the claim.
We start by proving the identifiability of the parameter for $K=2.$
@@
Let $K=2$ and let $d_{1},d_{2}>0.$ If the allele frequencies of $p_{\cdot,i}$ and $p_{\cdot,i+2}$ are linearly independent and in addition $p_{1,m}\neq p_{2,m}\forall m\in\{i,i+1,i+2\}$, the parameters $(q,r)$ are identifiable based on the random variables $X_{i},...,X_{i+2}.$
@@
and $M_{1},M_{2},M_{3}]$ is a three-dimensional array for which the element $(x_{1},x_{2},x_{3})$ equals $M_{1},M_{2},M_{3}]_{x_{1},x_{2},x_{3}}.$ Without loss of generality, we assume $i=1$. It holds
$\mathbb{P}(X_{1}$ $=x_{1},X_{2}=x_{2},X_{3}=x_{3})$
$=\sum_{z_{1},z_{2},z_{3}}\pi_{z_{1}}A_{1}(z_{1},z_{2})A_{2}(z_{2},z_{3})B_{2}(x_{2},z_{2})B_{3}(x_{3},z_{3})B_{1}(x_{1},z_{1})$
$=[\pi B_{1}A_{1},B_{2},\underbrace{A_{2}B_{3}}_{=:M_{2}}]_{x_{1},x_{2},x_{3}}.$
@@
    prove that the entries of $M_{1}:=B_{1}A_{1},M_{2},B_{2}$ are identifiable from the marginal density of $X_{1},X_{2},X_{3}$. Therefore, we use theory by Allman et al. (2009); Kruskal (1977).
@@
Hence, $B_{2}$ has the full rang if and only if $q_{1}\neq\frac{p_{1,2}}{p_{1,2}+p_{2,2}}.$ We have
@@
$a$ $=\mathbb{P}(X_{1}=0|Z_{2}=1)$
$=\sum_{z_{1}=1}^{2}\mathbb{P}(X_{0}=0|Z_{1}=z_{1})\mathbb{P}(Z_{1}=z_{1}|Z_{2}=1)$
$=(1-q_{1}p_{1,1})\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right)+\left(1-(1-q_{1})p_{2,1}\right)(1-e^{-rd_{1}})(1-q_{1})$
@@
$b$ $=\mathbb{P}(X_{1}=1|Z_{2}=1)$
$=q_{1}p_{1,1}\left((1-e^{-rd_{1}})q_{1}+e^{-rd_{1}}\right)+p_{2,1}(1-q_{1})^{2}(1-e^{-rd_{1}}),$
$c$ $=\mathbb{P}(X_{1}=0|Z_{2}=2)$
$=(1-q_{1}p_{1,1})(1-e^{-rd_{1}})(1-q_{1})+\left(1-(1-q_{1})p_{2,1}\right)\left((1-e^{-rd_{1}})(1-q_{1})+e^{-rd_{1}}\right),$
$d$ $=\mathbb{P}(X_{1}=1|Z_{2}=2)$
$=q_{1}p_{1,1}(1-e^{-rd_{1}})(1-q_{1})+p_{2,1}(1-q_{1})\left((1-e^{-rd_{1}})(1-q_{1})+e^{-rd_{1}}\right).$
@@
$\det(M_{1})=(2e^{-2rd_{1}}q_{1}^{2}-3e^{-2rd_{1}}q_{1}+e^{-2rd_{1}}-4e^{-rd_{1}}q_{1}^{2}+6e^{-rd_{1}}q_{1}-3e^{-rd_{1}}+2q_{1}^{2}-3q_{1}+1)$
$\quad\quad\quad\quad\quad\quad\cdot(p_{1,1}q_{1}+p_{2,1}q_{1}-p_{2,1})=0$
$\Leftrightarrow q_{1}=\frac{p_{2,1}}{p_{1,1}+p_{2,1}}$
$\quad\quad\lor(2e^{-2rd_{1}}q_{1}^{2}-3e^{-2rd_{1}}q_{1}+e^{-2rd_{1}}-4e^{-rd_{1}}q_{1}^{2}+6e^{-rd_{1}}q_{1}-3e^{-rd_{1}}+2q_{1}^{2}-3q_{1}+1)=0.$
It holds
$2e^{-2rd_{1}}q_{1}^{2}-3e^{-2rd_{1}}q_{1}+e^{-2rd_{1}}-4e^{-rd_{1}}q_{1}^{2}+6e^{-rd_{1}}q_{1}-3e^{-rd_{1}}+2q_{1}^{2}-3q_{1}+1=0$
$\Leftrightarrow q_{1}=\frac{e^{-rd_{1}}}{e^{-rd_{1}}-1}\lor q_{1}=\frac{e^{-rd_{1}}-3}{2(e^{-rd_{1}}-1)}.$
Fortunately, it holds $\frac{e^{-rd_{1}}}{e^{-rd_{1}}-1}\leq 0$ and $\frac{e^{-rd_{1}}-3}{2(e^{-rd_{1}}-1)}\geq 3/2.$ Hence, we do not have to take care of these terms.
We additionally calculate
$\mathbb{P}(X_{k-2}$ $=x\mid Z_{k-1}=z)$
$=\sum_{z^{\prime}=1}^{2}\mathbb{P}(X_{k-2}=x,Z_{k-2}=z^{\prime}\mid Z_{k-1}=z)$
$\propto\sum_{z^{\prime}=1}^{2}\mathbb{P}(X_{k-2}=x\mid Z_{k-2}=z^{\prime})\cdot\mathbb{P}(Z_{k-2}=z^{\prime}\mid Z_{k-1}=z)$
@@
$\propto(q_{1}p_{1,3})^{x}(1-q_{1}p_{1,3})^{1-x}\left(1_{z=1}e^{-rd_{2}}+q_{z}(1-e^{-rd})\right)$
$\hskip 28.45274pt+(q_{2}p_{2,3})^{x}(1-q_{2}p_{2,3})^{1-x}\left(1_{z=2}e^{-rd_{2}}+q_{z}(1-e^{-rd})\right).$
@@
$a_{2}$ $=(1-q_{1}p_{1,3})\left(e^{-rd_{2}}+q_{1}(1-e^{-rd_{2}})\right)+(1-(1-q_{1})p_{2,3})q_{1}(1-e^{-rd_{2}})$
$b_{2}$ $=(q_{1}p_{1,3})\left(e^{-rd_{2}}+q_{1}(1-e^{-rd_{2}})\right)+(1-q_{1})p_{2,3}q_{1}(1-e^{-rd_{2}}),$
$c_{2}$ $=(1-q_{1}p_{1,3})(1-q_{1})(1-e^{-rd_{2}})$
$\quad\quad+(1-(1-q_{1})p_{2,3})\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right),$
$d_{2}$ $=q_{1}p_{1,3}(1-q_{1})(1-e^{-rd_{2}})+(1-q_{1})p_{2,3}\left(e^{-rd_{2}}+(1-q_{1})(1-e^{-rd_{2}})\right).$
It holds
$\det(M_{2})=e^{-rd_{2}}(p_{1,3}q_{1}+p_{2,3}q_{1}-p_{2,3})=0$
$\Leftrightarrow r=\infty\lor q_{1}=\frac{p_{2,3}}{p_{1,3}+p_{2,3}}.$
For $r=\infty,$ it holds $\mathbb{P}(X_{1}=x)=(q_{1}p_{1,1}+(1-q_{1})p_{2,1})^{x}(1-(q_{1}p_{1,1}+(1-q_{1})p_{2,1}))^{1-x}.$ Then, $q$ is identifiable if and only if $p_{2,1}\neq p_{1,1}.$
Now, we come to step (ii), i.e. we have to ensure that the parameters are uniquely determined based on the entries. For the emission probabilities, this is a direct consequence of the assumption $p_{k,m}\neq p_{\ell,m}.$ Then, since $d_{m}>0,$ it is also immediate to infer the value $r$ from the other parameters.
@@
Now, we aim to infer (4) from the identifiability. Therefore, we first prove a law of large numbers. Then, based on this, we conclude (4).
To prove Theorem 4, we will use the following particularity of the hidden chain.
@@
Additionally, we will use that the hidden chain is uniformly ergodic (Lemma 5.4), which is defined as follows.
@@
Lemma 5.4 (Uniformly Ergodicity in the Linkage Model).
@@
According to Saloff-Coste and Zúniga (2007), Theorem 3.3, the convergence result holds from Lemma 5.2. ∎
@@
Assumption 3.1 implies that there exists a function $h_{i}$ and an integer $s\in\mathbb{N}$ such that
$(i)\quad\frac{1}{n}\sum_{i=1}^{n}h_{i}(X_{i},...,X_{i+s})$ $\xrightarrow[]{n\to\infty}1\quad\mathbb{P}_{(q^{0},r^{0})}\,-a.s.$
$(ii)\quad\frac{1}{n}\sum_{i=1}^{n}h_{i}(X_{i},...,X_{i+s})$ $\xrightarrow[]{n\to\infty}0\quad\mathbb{P}_{(q,r)}\,-a.s.$
@@
The proof is based on Lemma 6 in Douc et al. (2011). The only difference is that in our case $X$ is not a stationary sequence. We find an integer $s$ and a bounded functions $h_{i},i=1,2,$..., such that $\mathbb{E}_{\pi}(h_{i}(X_{i},...,X_{i+s}))=1,\mathbb{E}_{\pi^{0}}(h_{i}(X_{i},...,X_{i+s}))=0$ according to the theorem from Hahn-Banach for $\pi\neq\pi^{0}$. More precisely, we choose $h_{i}=1-\frac{f_{i}-1}{\mathbb{E}_{\pi^{0}}(f_{i})-1}$ with $f_{i}:=\frac{\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s})}{\mathbb{P}_{\pi}(X_{i},...,X_{i+s})}$ for $\theta\neq\theta^{0}$. The case $\mathbb{E}_{\pi^{0}}(f_{i})=1$ cannot occur according to Lemma 5.1. However, we have to ensure that there exists a constant $c_{q,r,q^{0},r^{0}}>0,$ such that $\mathbb{E}_{\pi^{0}}(f_{i})\geq 1+c_{q,r,q^{0},r^{0}}.$ It holds
$\mathbb{E}_{\pi^{0}}(f_{i})$ $=1+\int\frac{(\mathbb{P}_{\pi}(X_{i},...,X_{i+s})-\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s}))^{2}}{\mathbb{P}_{\pi}(X_{i},...,X_{i+s})}dX_{i}....X_{i+s}$
$\geq 1+\frac{1}{\kappa^{2(s+1)}}\int(\mathbb{P}_{\pi}(X_{i},...,X_{i+s})-\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s}))^{2}\,dX_{i}....X_{i+s}$
$\geq 1+\frac{1}{\kappa^{2(s+1)}K^{s+1}}|\mathbb{P}_{\pi}(X_{i},....,X_{i+s})-\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s})|_{TV}$
for $\kappa=\kappa_{q}\lor\kappa_{p}.$ There exists a $\delta>0$, such that $|\mathbb{P}_{\pi}(X_{m}=x|Z_{m}=z)-\mathbb{P}_{\pi^{0}}(X_{m}=x|Z_{m}=z)|\geq\delta$ for all $z_{m},z,x,z_{m-1}$, the total variation distance between $\mathbb{P}_{\pi^{0}}(X_{i},...,X_{i+s})$ and $\mathbb{P}_{\pi}(X_{i},...,X_{i+s})$ is bounded below.
The densities are uniformly bounded according to Assumption 3.1, i.e. $h_{i}$ is uniformly bounded.
Hence, we proceed similar to the proof of Theorem 3.12 in Dolgopyat and Sarig (2023) to prove (i) and (ii). Dolgopyat and Sarig (2023) considered the convergence of $\sum_{i=1}^{n}f_{i}(Z_{i},Z_{i+1})$ under certain conditions. Here, we consider the convergence of $\sum_{i=1}^{n}f_{i}(g_{i}(Z_{i}),g_{i+1}(Z_{i+1}))$ under certain conditions. The only difference compared to the proof in Dolgopyat and Sarig (2023) is that we have to prove
$\|\mathbb{E}(f_{i}(g_{i}(Z_{i}),g_{i+1}(Z_{i+1}))-\mathbb{E}(f_{i}(g_{i}(Z_{i}),g_{i+1}(Z_{i+1}))|Z_{j})\|_{2}\leq c\rho^{i-j}$
for $i>j.$ However, this claim is a direct consequence of Lemma 5.4.
We use Lemma 5.5 to prove the identifiably of the parameters. This is stated in the following theorem (Theorem 7.13 in van Handel (2008)).
@@
Therefore, we use Shiryaev (2016). Since the number of observable states is finite, the claim is a direct consequence of this.
Now, we prove that $\mathbb{P}_{q^{0},r^{0}}\neq\mathbb{P}_{q,r}$ implies that $\mathbb{P}_{q^{0},r^{0}},\mathbb{P}_{q,r}$ are mutually singular. This is Lemma 5.5. ∎